OpenStudy (anonymous):

i'm not to sure on this one, A bike roas race starts a an elevation of 870 feet and passes through five staes where the elevation changes by -684 feet, 17 feet, -236feet , -440feet and 286 feet. at what elevation does the race end?

OpenStudy (anonymous):

870-684+17+....

OpenStudy (anonymous):

translate this to a mathematical equation and solve

OpenStudy (anonymous):

so would teanswer be 285 feet

OpenStudy (anonymous):

you just do the arithmetic and add all the numbers together.

OpenStudy (anonymous):

isn't -187 the final elevation not the change.

OpenStudy (anonymous):

aaron is correct you add them as they occur 870 + (-684) + 17 + (-236) + (-440) + 286 = 870 -684+17-236-440+286 = -187 The race ends at an elevation of -187 feet

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

pls click GOOD ANSWER button ( :}))